# Dual Caption Preference Optimization for Diffusion Models

### v1

@@ 3 Experiments
Table 1: Results on PickScore, HSPv2.1, ImageReward (normalized), and CLIPScore. We show that DCPO significantly improves on Pickscore, HSP2.1, and ImageReward.
[…]
@@
We fine-tuned the U-Net model of Stable Diffusion 2.1 (SD 2.1) using DCPO on the Pick-Double Caption dataset and compared it to SD 2.1 models fine-tuned with $\text{SFT}_{\text{Chosen}}$, Diffusion-DPO, and MaPO on Pick-a-Pic v2 across various metrics. We first describe the Pick-Double Caption dataset and compare it to Pick-a-Pic v2. Subsequently, we provide an in-depth analysis of the results. More comparisons are found in Appendix C. Also, for further details on the fine-tuning, refer to Appendix F.
@@ 3.1 Pick-Double Caption Dataset
Table 2: Results on the GenEval Benchmark. DCPO significantly enhances model performance in generating the correct number of objects, improving image quality in terms of colors, and constructing attributes accurately.
[…]
@@
SD 2.1222Note that we rerun all the models on same seeds to have a fair comparison. 0.4775 0.96 0.52 0.35 0.80 0.09 0.15
@@
Motivated by the conflict distribution observed in previous preference datasets, we applied the captioning method described in Section 2.2.1 to generate unique captions for each image in the Pick-a-Pic v2 dataset. For the Pick-Double Caption dataset, we sampled 20,000 instances from Pick-a-Pic v2 and cleaned the samples as detailed in Appendix D. We then employed two state-of-the-art captioning models, LLaVa-1.6-34B and Emu2-37B, to generate captions for both the preferred and less preferred images, as illustrated in Figure 2.
To generate the captions, we used two different prompting strategies: 1) Conditional prompt: where the model was explicitly instructed to generate a caption for image $x$ based on the given prompt $c$, and 2) Non-conditional prompt: where the model provided a general description of the image in one sentence without referring to a specific prompt. More details are in Appendix D.
We evaluated the captions generated by LLaVA and Emu2 using CLIPscore, which revealed several key insights. LLaVA produced captions that have more correlation with the images for both preferred and less preferred samples compared to Emu2 and the original captions, although LLaVA’s captions were significantly longer (see Table 6 in Appendix D). Models fine-tuned on captions from the conditional prompt strategy outperformed those using the non-conditional approach, though the conditional prompt captions were twice as long. Interestingly, despite Emu2 generating much shorter captions, the models fine-tuned on Emu2 were comparable to those fine-tuned on the original prompts from Pick-a-Pic v2.
Finally, we observe that the key advantage of the Pick-Double Caption dataset is the greater difference in CLIPscore ($\Delta\mu$) between preferred and less preferred images compared to the original prompts. Specifically, while the original prompt has a $\Delta\mu$ of 1.3, LLaVA shows a much larger difference at 4.3, and Emu2 at 2.8. This increased gap reflects improved alignment performance in models fine-tuned on this dataset, indicating that the captioning method mitigates the conflict distribution.
@@ 3.2 Performance Comparisons
We evaluated all methods on 2,500 unique prompts from the Pick-a-Pic v2 (Kirstain et al., 2023) dataset, measuring performance using Pickscore (Kirstain et al., 2023), CLIPscore (Hessel et al., 2022), and Normalized ImageReward (Xu et al., 2023). Additionally, we generated images from 3,200 prompts in the HPSv2 (Wu et al., 2023b) benchmark and evaluated them using the HPSv2.1 model. To provide a comprehensive evaluation, we also compared the methods using GenEval (Ghosh et al., 2023), focusing on how well the fine-tuned models generated images with the correct number of objects, accurate colors, and proper object positioning.
We compared different versions of DCPO, including the captioning (DCPO-c), perturbation (DCPO-p), and hybrid (DCPO-h) methods, with other approaches, as outlined in Section 2.2. For more information on the fine-tuning process of the models, refer to Appendix F.
The results in Tables 1 and 2 show that DCPO-h significantly outperforms the best scores from other methods, with improvements of +0.21 in Pickscore, +0.45 in HPSv2.1, +1.8 in ImageReward, +0.15 in CLIPscore, and +3% in GenEval. Additionally, the results demonstrate that DCPO-c outperforms all other methods on GenEval, Pickscore, and CLIPscore. While DCPO-p performs slightly worse than DCPO-c, it still exceeds SD 2.1, SFT, Diffusion-DPO, and MaPO on GenEval. However, its scores on ImageReward and Pickscore suggest that it underperforms compared to the other approaches. Importantly, DCPO-p shows significant improvement over the other methods on HPSv2.1, highlighting the effectiveness of the perturbation method.
3.3 Ablation Studies and Analysis
Support of Hypothesis 1.
Support of Hypothesis 2.
To illustrate the impact of the correlation between the prompt $c$ and image $x$ on the perturbation method, we perturbed both the original prompt $c$ and the less preferred caption $z^{w}$, generated by the model $Q_{\phi}$, where $z^{w}\sim W_{\phi}(z^{w}|Q_{\phi}(z^{w}|x^{w},c))$. At the same time, we kept the caption generated by $Q_{\phi}$ for the preferred image as the preferred caption, $z^{w}\sim Q(z^{w}|x^{w},c)$. In this case, we assume $Q_{\phi}=\text{LLaVA}$ and $W_{\phi}=\text{DIPPER}$. The results in Table 6 in Appendix D show that the caption $z$ generated by LLaVA is more correlated with the image $x$ than the original prompt $c$, indicating that $S(z,x)>S(c,x)$. Based on the results in Table 3, we conclude that perturbing more correlated captions leads to better performance.
@@ 3.3 Ablation Studies and Analysis
In- vs. Out-of Distribution.
Inspired by (Verma et al., 2024), which examined the impact of out-of-distribution data in vision language models, we assessed DCPO using both in-distribution and out-of-distribution (OOD) data. As discussed in Section 3.1, the captioning model can generate OOD captions. To explore this, we fine-tuned SD 2.1 with DCPO-h using LLaVA and Emu2 captions at a medium perturbation level. Figure 6 shows that in-distribution data improve alignment performance, while OOD results for LLaVA in GenEval, Pickscore, and CLIPscore are comparable to Diffusion-DPO. Similar behavior was observed for DCPO-c, as noted in Appendix F.
@@
DCPO-h (LLaVA) Pair ($z^{w}$,$z^{l}_{p}$) (32.32, 31.17) 20.57 25.62 58.2 27.13 0.5100
DCPO-h (LLaVA) Pair ($z^{w}$,$z^{w}_{p}$) (32.32, 27.01) 20.57 25.27 57.4 27.18 0.5110
@@
Figure 7: DCPO-h performance comparison across various $\beta$ values, evaluated on HPSv2.1 and GenEval.
Explore on $\beta$.
In DCPO, $\beta$ is a key hyperparameter. To evaluate its impact, we fine-tuned SD 2.1 using different values of $\beta=\{500,1000,1500,2500,5000\}$. Interestingly, in Figure 7 we observed that $\beta=500$ showed significant improvements on HPSv2.1 and GenEval, even surpassing DCPO-h with $\beta=5000$, our best-reported model. Additional results for different $\beta$ values can be found in Appendix F.
@@
We evaluated DCPO-h and Diffusion-DPO using GPT-4o on the PartiPrompts benchmark, consisting of 1,632 prompts. GPT-4o assessed images based on three criteria: Q1) General Preference (Which image do you prefer given the prompt?), Q2) Visual Appeal (Which image is more visually appealing?), and Q3) Prompt Alignment (Which image better fits the text description?). As shown in Figure 8, DCPO-h outperformed Diffusion-DPO in Q1 and Q2, with win rates of 58% and 66%. To see the style of the prompts, refer to Appendix G.
@@ 4 Related Works
Recent advances in preference alignment methods of text-to-image diffusion models have shown that RL-free methods (Wallace et al., 2024; Yang et al., 2024; Li et al., 2024; Yuan et al., 2024; Gambashidze et al., 2024; Park et al., 2024) outperforms RL-based approaches (Fan & Lee, 2023; Fan et al., 2023; Hao et al., 2023; Lee et al., 2023; Xu et al., 2023; Prabhudesai et al., 2024; Black et al., 2024; Clark et al., 2024) mainly because they eliminates the need for an explicit reward model (Saeidi et al., 2024a; Chatterjee et al., 2024a). Initially, Direct Preference Optimization (DPO) (Rafailov et al., 2024) and Triple Preference Optimization (Saeidi et al., 2024b) method reformulate the RLHF objective in a closed-form manner and introduce it as an implicit reward model with a simple classification objective. Diffusion-DPO (Wallace et al., 2024) directly adopts DPO method into text-to-image diffusion models, utilizing pairwise preference datasets consisting of text and images to guide alignment. Diffusion-KTO (Li et al., 2024) incorporates Kahneman & Tversky model of human utility to align these models, simplifying the process by using images with binary feedback signals, i.e., likes or dislikes instead of pairwise preference data. To enhance flexibility, Hong et al. (2024) introduce MaPO, an alignment technique independent of a reference model previously used by other methods, enabling greater control over stylistic adaptations. However, previous methods optimize diffusion models based on a single prompt for a pair of images, which supports the irrelevant prompts issue explored in Section 2.1.
@@
Text-to-image image preference datasets commonly involve the text prompt to generate the images, and two or more images are ranked according to human preference. HPS (Wu et al., 2023c) and HPSv2 (Wu et al., 2023b) create multiple images using a series of image generation models for a single prompt, and the images are ranked according to real-world human preferences. Moreover, a classifier is trained using the gathered preference dataset, which can be used as a metric for image-aligning tasks. Also, Pick-a-Pic v2 (Kirstain et al., 2023) follows a similar structure to create a pairwise preference dataset along with their CLIP (Radford et al., 2021) based scoring function, Pickscore. While these datasets are carefully created, having only one prompt for both or all the images introduces conflict distribution, which will be further discussed in Section 2.1. For this reason, we modified the Pick-a-Pic v2 dataset using recaptioning and perturbation methods to improve image alignment performance.
5 Conclusion and Limitations
In this paper, we present a novel preference optimization method for aligning text-to-image diffusion models called Dual Caption Preference Optimization (DCPO). We tackle two major challenges in previous preference datasets and optimization algorithms: the conflict distribution and irrelevant prompt. To overcome these issues, we introduce the Pick-Double Caption dataset, a modified version of the Pick-a-Pic v2 dataset. We also identify difficulties in generating captions, particularly the risk of out-of-distribution captions for images, and propose three approaches: 1) captioning (DCPO-c), 2) perturbation (DCPO-p), and 3) a hybrid method (DCPO-h). Our results show that DCPO-h significantly enhances alignment performance, outperforming methods like MaPO and Diffusion-DPO across multiple metrics.
Limitations.
Although DCPO shows strong performance across various metrics, the captioning and perturbation methods are resource-intensive. We encourage future research to explore cost-effective alternatives to these methods. Additionally, the potential of using different backbones, such as Stable Diffusion XL (SDXL) (Rombach et al., 2022), has not been explored in the context of DCPO. We also invite researchers to investigate DCPO’s effectiveness on other tasks, such as safety. We believe our work will have a significant impact on the alignment research community.

### v2

@@ 4 Experiments
Table 1: Results on PickScore, HPSv2.1, ImageReward (normalized), and CLIPScore. We show that DCPO significantly improves on Pickscore, HPSv2.1, and ImageReward.
    Pick  Score   ($\uparrow$) HPSv2.1 ($\uparrow$)       Image  Reward     ($\uparrow$)     CLIP  Score   ($\uparrow$)
@@
We fine-tuned the U-Net model of Stable Diffusion (SD) 2.1 using DCPO on the Pick-Double Caption dataset and compared it with SD 2.1 models fine-tuned with $\text{SFT}_{\text{Chosen}}$, Diffusion-DPO, and MaPO on Pick-a-Picv2 in various metrics. We first describe the Pick-Double Caption dataset and compare it to Pick-a-Picv2. Subsequently, we provide an in-depth analysis of the results. Details on fine-tuning are in Appendix F, and further comparisons are in Appendix B.
@@ 4.1 Pick-Double Caption Dataset
Motivated by the semantic overlap observed in previous preference datasets, we applied the captioning method described in Section 3.2.1 to generate unique captions for each image in the Pick-a-Pic v2 dataset. For the Pick-Double Caption dataset, we sampled 20,000 instances from Pick-a-Pic v2 and cleaned the samples as detailed in Appendix C. We then employed two state-of-the-art captioning models, LLaVa-1.6-34B and Emu2-37B, to generate captions for both the preferred and less preferred images, as shown in Figure 2.
To generate the captions, we used two different prompting strategies: 1) Conditional prompt: where the model was explicitly instructed to generate a caption for image $x$ based on the given prompt $c$, and 2) Non-conditional prompt: where the model provided a general description of the image in one sentence without referring to a specific prompt. More details are in Appendix C.
We evaluated the captions generated by LLaVA and Emu2 using CLIPscore, which revealed several key insights. LLaVA produced captions that have more correlation with the images for both preferred and less preferred samples compared to Emu2 and the original captions, although LLaVA’s captions were significantly longer (see Table LABEL:tab:tokens-detail in Appendix C). Models fine-tuned on captions from the conditional prompt strategy outperformed those using the non-conditional approach, though the conditional prompt captions were twice as long. Interestingly, despite Emu2 generating much shorter captions, the models fine-tuned on Emu2 were comparable to those fine-tuned on the original prompts from Pick-a-Pic v2.
Finally, we observe that the key advantage of the Pick-Double Caption dataset is the greater difference in CLIPscore ($\Delta\mu$) between preferred and less preferred images compared to the original prompts. Specifically, while the original prompt has a $\Delta\mu$ of 1.3, LLaVA shows a much larger difference at 4.3, and Emu2 at 2.8. This increased gap reflects improved alignment performance in models fine-tuned on this dataset, indicating that the captioning method mitigates the semantic overlap.
Table 2: Results on the GenEval Benchmark. DCPO contributes to model performance in generating the correct number of objects, improving image quality in terms of colors, and constructing attributes accurately.
Method Overall     Single  object     Twoobjects Counting Colors Position       Attribute   binding
@@
SD 2.1222Note that we rerun all the models on the same seeds to have a fair comparison. 0.4775 0.96 0.52 0.35 0.80 0.09 0.15
@@ 4.2 Performance Comparisons
As shown in Table 1, we evaluated all methods on 2,500 unique prompts from the Pick-a-Picv2 (Kirstain et al., 2023) dataset, measuring performance using Pickscore (Kirstain et al., 2023), CLIPscore (Hessel et al., 2022), and Normalized ImageReward (Xu et al., 2023). We also generated images from 3,200 prompts in the HPSv2 (Wu et al., 2023b) benchmark and evaluated them using the HPSv2.1 model. To provide a more fine-grained evaluation, in Table 2, we also compared the methods using GenEval (Ghosh et al., 2023), focusing on how well the fine-tuned models generated images with the correct number of objects, accurate colors, and proper object positioning.
We compared our different variants of DCPO, including the captioning (DCPO-c), perturbation (DCPO-p), and hybrid (DCPO-h) methods, with other approaches, as outlined in Section 3.2. For more information on the fine-tuning process of the models, refer to Appendix F.
@@
The results in Tables 1 and 2 show that DCPO-h significantly outperforms the best scores from other methods, with improvements of +0.21 in Pickscore, +0.31 in HPSv2.1, +1.8 in ImageReward, +0.15 in CLIPscore, and +2% in GenEval. Additionally, the results demonstrate that DCPO-c outperforms all other methods on GenEval, Pickscore, and CLIPscore. While DCPO-p performs slightly worse than DCPO-c, it still exceeds SD 2.1, SFT, Diffusion-DPO, and MaPO on GenEval. However, its scores on ImageReward and Pickscore suggest that it underperforms compared to the other approaches. Importantly, DCPO-p shows significant improvement over the other methods on HPSv2.1, highlighting the effectiveness of the perturbation method.
4.3 Ablation Studies and Analyses
Support of Hypothesis 1.
Support of Hypothesis 2.
To illustrate the impact of the correlation between the prompt $c$ and image $x$ on the perturbation method, we perturbed both the original prompt $c$ and the less preferred caption $z^{w}$, generated by the model $Q_{\phi}$, where $z^{w}\sim W_{\phi}(z^{w}|Q_{\phi}(z^{w}|x^{w},c))$. At the same time, we kept the caption generated by $Q_{\phi}$ for the preferred image as the preferred caption, $z^{w}\sim Q(z^{w}|x^{w},c)$. In this case, we assume $Q_{\phi}=\text{LLaVA}$ and $W_{\phi}=\text{DIPPER}$. The results in Table LABEL:tab:tokens-detail in Appendix C show that the caption $z$ generated by LLaVA is more correlated with the image $x$ than the original prompt $c$, indicating that $S(z,x)>S(c,x)$. Based on the results in Table 3, we conclude that perturbing more correlated captions leads to better performance.
@@ 4.3 Ablation Studies and Analyses
DCPO-h (LLaVA) Pair ($z^{w}$$z^{l}_{p}$) (32.32, 31.17) 20.57 25.62 58.2 27.13 0.5100
DCPO-h (LLaVA) Pair ($z^{w}$$z^{w}_{p}$) (32.32, 27.01) 20.57 25.27 57.4 27.18 0.5110
In- vs. Out-of Distribution.
We evaluated DCPO on in-distribution and out-of-distribution (OOD) data. As discussed in Section 4.1, the captioning model can generate OOD captions. To explore this, we fine-tuned SD 2.1 with DCPO-h using LLaVA and Emu2 captions at a medium perturbation level. Figure 6 shows that in-distribution data significantly improve alignment performance, while OOD results for LLaVA in GenEval, Pickscore, and CLIPscore are comparable to Diffusion-DPO. Similar behavior was observed for DCPO-c, as noted in Appendix F.
@@
Hyperparameter Tuning.
We conducted a systematic hyperparameter study for a fair comparison across DCPO and all baselines. Under a controlled protocol with batch size fixed at $128$ and learning rate fixed at $(1\times 10^{-8})$, we swept $\beta$ for DCPO and Diffusion-DPO over the set $[500,1000,2000,2500,5000]$. DCPO achieved its strongest performance at a $\beta$ value of $5000$, outperforming alternatives on Pickscore, GenEval, HPSv2.1, ImageReward, and CLIPscore, while Diffusion-DPO reached its optimum at a $\beta$ value of $2000$. SFT performed best with batch size $64$ and learning rate $(1\times 10^{-9})$. MaPO attained its optimum at a much smaller coefficient, with a $\beta$ value of $0.01$, indicating a preference for lighter regularization relative to DCPO and DPO. These outcomes align with the objectives of each method, where MaPO’s margin-oriented formulation is sensitive to over-weighting the margin and thus favors small $\beta$, DCPO’s diffusion-consistent objective with dual captions scales the preference signal while preserving stable updates and thus benefits from larger $\beta$, and Diffusion-DPO’s balance between the preference term and a reference-model KL naturally favors a moderate $\beta$. For additional details on the hyperparameter tuning process refer to Appendix F.
@@
We evaluated DCPO-h and Diffusion-DPO using GPT-4o on the PartiPrompts benchmark, consisting of 1,632 prompts. GPT-4o assessed images based on three criteria: Q1) General Preference (Which image do you prefer given the prompt?), Q2) Visual Appeal (Which image is more visually appealing?), and Q3) Prompt Alignment (Which image better fits the text description?). As shown in Figure 7, DCPO-h outperformed Diffusion-DPO in Q1 and Q2, with win rates of 58% and 66%. Beyond GPT-4o, we conducted a small human study for Q2 to measure the alignment of GPT-4o with human preference; results in Table 11 in Appendix G indicate that GPT-4o’s judgments align closely with human preferences. See Appendix G for more details.
Figure 8: Benchmarking results by alternative perturbation selection mechanisms (MLLM-as-a-Judge models) alongside our original CLIPScore-based judge. All experiments are under the same SD 2.1 and DCPO-h + LLaVa captions setting. For clarity, we only show the averages across all perturbation levels, per perturbation judge and benchmark.
Alternative MLLM-as-a-Judge for Controllable Perturbation.
We use CLIPScore as the selection mechanism for incorporating the three levels of perturbations used in our PickDouble dataset. To confirm the effectiveness of CLIPScore-as-a-Judge on perturbation selection, we further conducted a series of controlled ablation studies using perturbations ranked by alternative MLLM-based judges inspired by Chen et al. (2024). We experimented with the following judge models - Qwen Chat, Qwen2.5-VL-7B, LlaVa1.6-7B, and LlaVa1.6-13B. The results in Figure 8 show that all MLLM-as-a-Judge variants consistently underperform compared to our CLIPScore-based baselines across all five downstream benchmarks, regardless of the judge model’s scale or the granularity of its perturbation strengths. This further validates that our original method identifies more effective perturbations (i.e., modified prompts) that lead to improved downstream performance across multiple benchmarks. More details can be found in Appendix E.
DCPO vs Online and Iterative Methods.
Although DCPO is an offline method, we compare it against online and iterative approaches such as SePPO Zhang et al. (2024) and SPO Liang et al. (2025) to assess effectiveness. Under a small budget of optimization steps and limited training data, DCPO consistently outperforms both, demonstrating superior sample and compute efficiency. By contrast, SPO typically requires substantially more training iterations and data to reach similar performance, and SePPO depends on iterative online updates to approach comparable results. These findings indicate that DCPO achieves strong alignment with fewer resources while avoiding the complexity of online optimization. Refer to Appendix B.3 for more details.
@@ 5 Related Works
Recent advances in preference alignment for text-to-image diffusion models show that RL-free methods (Yang et al., 2024a; Li et al., 2024; Yuan et al., 2024; Gambashidze et al., 2024; Park et al., 2024) outperform RL-based ones (Fan & Lee, 2023; Fan et al., 2023; Hao et al., 2023; Lee et al., 2023; Prabhudesai et al., 2024; Black et al., 2024; Clark et al., 2024) by removing the need for explicit reward models. Methods like Diffusion-DPO (Wallace et al., 2024), which adapts DPO (Rafailov et al., 2024), and Diffusion-KTO (Li et al., 2024), which uses binary feedback instead of pairwise data, streamline alignment. MaPO (Hong et al., 2024) further increases flexibility by eliminating the dependency on reference models. However, these methods often align based on single-prompt image pairs, leading to issues with irrelevant prompts (see Section 3.1). While online preference optimization shows strong performance (Fan et al., 2023; Yang et al., 2024b; Lou et al., 2024), it requires a reward model and more optimization steps, making it vulnerable to reward hacking. In this work, we focus on comparing DCPO with offline methods that do not rely on extra models for ranking of diffusion outputs.
@@
Text-to-image image preference datasets commonly involve the text prompt to generate the images, and two or more images are ranked according to human preference. HPS (Wu et al., 2023c) and HPSv2 (Wu et al., 2023b) create multiple images using a series of image generation models for a single prompt, and the images are ranked according to real-world human preferences. Moreover, a classifier is trained using the gathered preference dataset, which can be used as a metric for image-aligning tasks. Also, Pick-a-Pic v2 (Kirstain et al., 2023) follows a similar structure to create a pairwise preference dataset along with their CLIP (Radford et al., 2021) based scoring function, Pickscore. While these datasets are carefully created, having only one prompt for both or all the images introduces semantic overlap, which will be further discussed in Section 3.1. For this reason, we modified the Pick-a-Pic v2 dataset using recaptioning and perturbation methods to improve image alignment performance.
6 Conclusion
In this paper, we present a novel preference optimization method for aligning text-to-image diffusion models called Dual Caption Preference Optimization (DCPO). We tackle two major challenges in previous preference datasets and optimization algorithms: the semantic overlap and irrelevant prompt. To overcome these issues, we introduce the Pick-Double Caption dataset, a modified version of the Pick-a-Pic v2 dataset. We also identify difficulties in generating captions, particularly the risk of out-of-distribution captions for images, and propose three approaches: 1) captioning (DCPO-c), 2) perturbation (DCPO-p), and 3) a hybrid method (DCPO-h). Our results show that DCPO-h significantly enhances alignment performance, outperforming methods like MaPO and Diffusion-DPO across multiple metrics.
Limitation
While DCPO achieves strong performance on various benchmarking metrics, its captioning and perturbation processes require considerable computational resources. We encourage future research to explore more efficient and cost-effective alternatives to further enhance its practicality. The preliminary results in Appendix B.5 indicate that DCPO outperforms Diffusion-DPO across different backbones, such as Stable Diffusion XL (SDXL) (Podell et al., 2023). However, further investigation into its performance with emerging state-of-the-art models remains essential. Additionally, exploring DCPO’s applications in safety-related tasks could be a valuable direction for future work. We believe our research will contribute significantly to the alignment community and inspire further advancements in this field.
Ethical Considerations
The authors state that in this work, AI assistants, specifically Grammarly and ChatGPT, were utilized to correct grammatical errors and restructure sentences.